OpenStudy (anonymous):

three consecutive terms of an aritmetic sequence have a sum of 12 and a product of -80. Find the three terms

OpenStudy (kc_kennylau):

Let the first term be \(a\) and the common difference be \(d\). Try creating 2 equations from that :)

OpenStudy (kc_kennylau):

Hint: the second time is \(a+d\) :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Let x=first term and k=common difference; (x)+(x+k)+(x+2k)=12 \\since it has been said that the sum is 12. 3x+3k=12 \\combining like terms x+k=4 \\multiplying both sides by 1/3 k=4-x \\this is the value of the common difference substitute: (x)(x+k)(x+2k)=-80 \\since it has been said that the product is -80 (x)[(x)+(4-x)][(x)+2(4-x)]=-80 \\substituting the value of k (x)(4)(8-x)=-80 \\performing the operation (x)(8-x)=-20 \\multiplying both sides by 1/4 x^2-8x-20=0 \\equating to zero gives (x-10)(x+2)=0 \\the two factors of the equation therefore: If x=10 and k=-6, your sequence will be; {10, 4, -2} If x=-2 and k=6, your sequence will be; {-2, 4, 10} That's the answer. :)